A production lot of size 100 is known to be 55% defective. A random sample of 12

items is selected without replacement. Determine the probability that there will be

no defective in the sample.


1
Expert's answer
2021-08-11T12:13:38-0400

"P=\frac{45}{100}\frac{44}{99}\frac{43}{98}...\frac{34}{89}\approx0.00002738."